OpenStudy (anonymous):

find the exact value for 2Cos45xSin45 , I dont get how it equals 1

jimthompson5910 (jim_thompson5910):

use the identity 2*sin(t)*cos(t) = sin(2t) So this means 2*sin(t)*cos(t) = sin(2t) 2*sin(45)*cos(45) = sin(2*45) 2*sin(45)*cos(45) = sin(90) 2*sin(45)*cos(45) = 1

OpenStudy (anonymous):

But I understand that it would be 2(sqrt2/2) x (sqrt2/2) = 2(sqrt4/4) = 2(2/4) but I dont know how it equals 1 after that

OpenStudy (anonymous):

okay nevermind I'm just stupid. thank you so much !

jimthompson5910 (jim_thompson5910):

no you're not, but i'm glad you got it